Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 596     Entry time: 17 Mar 2025, 09:36
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 17.03.2025 
General Announcement:
- As DRC, Jola S.-D. says it was a quiet weekend.	 

CTRL:
- Nothing to report.

DA:
- Not present.

EEE:
- Nothing to report.

VAC:
- Nothing to report.

XPD:
- Nothing to report.
DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- Not present.

LAS:
- Not present.

UNSYS:
- Nothing to report; all fine.

SPB:
- Nothing to report; all good downstairs.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- All fine at HED;
- Question: possibility to use beam parasitically tomorrow? A. Galler says one can always open the shutter if stability or tuning is not required. 
Please contact the PRC for details on status or help with communication with BKR.

SCS:
- Not present.

SQS:
- Nothing to report.

SXP:
- Preparing for this week's experiment.
ELOG V3.1.4-7c3fd00